Soybean meal is 12% protein, cornmeal is 6% protein. How many pounds of each should be mixed together in order to get a 240 lb mixture that is 9% protein?

Answer by nerdybill(5411)

Soybean meal is 12% protein, cornmeal is 6% protein. How many pounds of each should be mixed together in order to get a 240 lb mixture that is 9% protein?
.
Let x = lbs of soybean meal in mixture
then because there is a 240 lb mixture we have
240-x = lbs of cornmeal in mixture
.
.12x + .06(240-x) = .09(240)
.12x + 14.4 - .06x = 21.6
.06x + 14.4 = 21.6
.06x = 7.2
x = 7.2/.06
x = 120 pounds of soybean meal
.
Cornmeal:
240-x = 240-120 = 120 pounds 0f cornmeal